OpenStudy (anonymous):

hey did you know how to solve (5-3/x)/(1+4/x^2)

OpenStudy (capture):

do you want to simplify? or is there an equality missing from your problem?

OpenStudy (anonymous):

simplify

OpenStudy (capture):

(5 - (3/x)) / (1+ (4/x^2)) = ( (5x-3)/3) / ( (x^2 +4) /x^2) = ( (5x-3)/3 ) * (x^2 / (x^2 + 4)) = ( 5x^2 - 3x) / (x^2 + 4)

OpenStudy (capture):

you can factor out an x as well if you want

OpenStudy (anonymous):

how?

OpenStudy (capture):

= ( x ( 5x - 3) ) / (x^2 + 4)

OpenStudy (anonymous):

thanks so much